OpenStudy (anonymous):

The value of [(x-y)^3 + (y-z)^3 + (z-x)^3]/[(9(x-y)(y-z)(z-x)] is equal to: a.) 1 b.) 0 c.) 1/3 d.) 1/9

OpenStudy (anonymous):

let x-y =a y-z =b z-x = c now a+b+c =0 hence you can proceed

OpenStudy (anonymous):

do you know the identity if a+b+c =0 a^3 + b^3 + c^3 = 3abc

OpenStudy (anonymous):

I was only familiar with square of binomials, stumbled upon that identity for the first time...

OpenStudy (anonymous):

if x-y =a y-z =b z-x = c then we have a+b+c=0 now the given question translates to (a^3 + b^3 + c^3 ) / 9abc = 3abc/9abc =1/3

OpenStudy (anonymous):

That helped, I just solved the problem... Thanks a lot! I have more follow up questions..

OpenStudy (anonymous):

need not mention..